rj mother iss 22 years older than febby and the sum of their ages 60. Find their ages
Answer by macston(5194)

m=mother=f+22; f=febby
m+f=60yrs Substitute for m.
f+22yrs+f=60yrs Subtract 22 from each side.
2f=38yrs Divide each side by 2.
f=19yrs ANSWER 1: Febby is 19 years old.
m=f+22=19+22=41yrs ANSWER 2: Mother is 41 years old.